# Liouville phenomenon for the Klein-Gordon equation in $1+1$ dimensions

### 04 Mar 2026 v1

@@ Abstract.
We study the Klein-Gordon equation in one spatial and one temporal dimension. Physically, this equation describes the wave function of a relativistic spinless boson with positive rest mass. Mathematically, this is the most elementary hyperbolic partial differential equation, after the wave equation itself. Relative to the point at the origin, the spacetime splits according to the light cones, and we find four quarter-planes, two of which are timelike while the remaining two are spacelike. Not unexpectedly, the solutions behave quite differently in the two types of quarter-planes. It turns out that the spacelike quarterplanes exhibit a Liouville phenomenon, where insufficient growth forces the solution to display a certain kind of symmetry, where the values along the two linear edges are in a one-to-one relation. This phenomenon shares features with the classical Liouville theorem in the theory of entire functions as well as the Phragmén-Lindelöf principle for harmonic functions.
@@ 1.1. Classical backgrund
The Klein-Gordon equation describes the wavefunction of a relativistic spinless boson with positive rest mass. In one spacial and one temporal dimension, it is advantageous to swith to using the characteristic directions to define alternative coaordinates, which after rescaling gives us the equation
@@
since $f(0)=g(0)$, $J_{0,0}(x,0)=1$, and $J_{1,0}(0,x-s)=0$. By interchanging the roles of $x$ and $y$ in the calculation, we find likewise that
@@
To offer a cohesive reader-friendly presentation, we supply the proof of this theorem in the appendix (Section 7). Unfortunately, we failed to pinpoint the exact reference to Riemann’s contribution here in his collected works [23] (perhaps the paper Ueber die Fortplantzung ebener Luftwellen von endlicher Schwingungsweite is the most relevant), but it is generally accepted that the method, which is mainly concerned with the Darboux-Goursat problem for inhomogeneous equations, is due to Riemann, and thus called the Riemann method.
@@
If we apply (1.6) to $a^{\prime}=x,b^{\prime}=y,a=b=0$, we may view the equation as a way to represent $u(x,y)$ in terms of $u(x,0)$, $u(0,y)$, $u(0,0)$, and the integral over the corresponding rectangle:
@@
Based on this quadrature identity, the Picard method applies to give uniqueness.
@@ 1.5. Splitting Riemann’s solution into unilateral wave solutions
This assertion is immediate from Theorems 1.2 and by slightly rearranging the terms in Riemann’s solution. Alternatively, we can observe that the expressed solution $u$ solves the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ in the sense of distribution theory since each of the terms $u_{0},u_{\mathrm{horz}},u_{\mathrm{vert}}$ does, and that the values along the the $x$-axis and the $y$-axis coincide with $f$ and $g$, respectively. Theorem 1.3 then shows that $u=u_{0}+u_{\mathrm{horz}}+u_{\mathrm{vert}}$ must coincide with $\mathbf{R}[f,g]$. Proposition 1.8 motivates the following terminology.
@@
It is the aim of this paper to study the growth properties of the unilateral wave solutions $u_{\mathrm{horz}}$ and $u_{\mathrm{vert}}$. By Lorentz group invariance, it is equivalent to focus on just one of them, and we will consider the horizontal unilateral wave solution $u_{\mathrm{horz}}$. As it turns out, there is Liouville phenomenon for such solutions. Here, we recall the classical theorem of Liouville, which asserts that any entire function which is also bounded has to be constant. There are many extensions, for instance, a harmonic function in the entire plane which is bounded must be constant as well. A version which is particularly attractive is known as the Phragmén-Lindelöf principle which asserts that if a harmonic function vanishes on a line and has insufficient growth on one side of the line, then it must vanish everywhere.
@@ 1.6. Motivation: the Phragmén-Lindelöf principle for harmonic functions
The connection with Liouville’s theorem is actually rather direct: By Schwarz reflection, which applies along the linear boundary since the values of the harmonic function vanish there, the harmonic function $u$ extends to a harmonic function in the whole complex plane with $u(\bar{z})=-u(z)$. The growth assumptions on this entire harmonic function are such that we have very slow growth. In other words, we are close in spirit to the classical Liouville theorem.
@@
There is a rather sizeable literature on Liouville type phenomena for elliptic equations, linear and nonlinear, also involving fractional Laplacians, etc. We could mention a particularly interesting effect in the discrete context, where harmonicity combined with boundedness on a big enough portion is enough to force the function to be constant [8]. This is in contrast with the continuous setting where it is enough to have a narrow path tending to infinity to permit the existence of a nontrivial harmonic function which grows wildly along the path while it is bounded elsewhere.
@@ 1.7. Liouville type phenomenon for spacelike quarter-planes
Apparently, in the context of the Phragmén-Lindelöf principle, there is some degree of minimal growth which must be surpassed to get a nontrivial solution. What about the unilateral solutions to the Klein-Gordon equation? By the characterization of Proposition 1.10, they vanish either on the horizontal line $\mathbb{R}\times\{0\}$ or the vertical line $\{0\}\times\mathbb{R}$. This may be able to force them to grow somewhere, and, in fact, as it turns out, the required growth is rather strong. A first result of this kind which we may present is the following. Note that with respect to the origin $(0,0)$, the equation $xy=0$ represents the light cone emanating from $(0,0)$, while $xy>0$ represents the timelike vectors, and $xy<0$ represents the spacelike vectors. The spacelike vectors can all be thought of as representing events contemporary with that of the origin.
@@
The proof of Theorem 1.14 is supplied in Subsection 2.3.
@@
(c) In the instance that $\mathcal{Y}=\mathbb{R}_{\leq 0}$, it appears that Theorem 1.14 was first obtained by Andrew Bakan (unpublished). After all, he mentioned a version of it in a seminar presentation in Stockholm during his visit in the fall semester of 2016. We have since learned that Bakan’s interest in the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ has its roots in the 1990s.
The condition that $u$ vanishes along the semi-axis $\mathbb{R}_{\geq 0}\times\{0\}$ entails that $u$ may be extended to a solution to the Klein-Gordon equation in the whole plane $\mathbb{R}^{2}$ which vanishes on the line $\{0\}\times\mathbb{R}$, that is, the extension is a horizontal unilateral solution.
@@
Suppose $\theta_{1},\theta_{2}\in\mathbb{R}_{>0}$ with $\theta_{1}\theta_{2}\geq 1$. There exists a nontrivial continuous horizontal unilateral solution $u$ to the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ such that the restriction of $u$ to the quarter-plane $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$ does not vanish identically, while it meets the uniform growth condition
@@
$\theta_{1}^{1/q}\theta_{2}^{1/q^{\prime\prime}}<\bigg(\frac{1}{q}\bigg)^{1/q}\bigg(\frac{1}{q^{\prime\prime}}\bigg)^{1/q^{\prime\prime}}\cos^{2}\frac{\pi(1-q)}{2q},$
@@
$\theta_{1}^{1/q}\theta_{2}^{1/q^{\prime\prime}}>\bigg(\frac{1}{q}\bigg)^{1/q}\bigg(\frac{1}{q^{\prime\prime}}\bigg)^{1/q^{\prime\prime}}\cos^{2}\frac{\pi(1-q)}{2q},$
@@ 2.1. An example of growth in a spacelike quadrant
holds uniformly. Moreover, since $u_{1}(x,0)=f_{1}(x)$ holds for $x\in\mathbb{R}$, it follows that $u_{1}$ is the nontrivial solution to the Klein-Gordon equation which is in particular does not vanish identically on $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$ and enjoys the claimed growth estimate. ∎
@@ 2.2. The evolution equation for horizontal unilateral wave solutions
(2.6) $\mathbf{L}[u_{\mathrm{horz}}(\cdot,y)](\zeta)=\int_{0}^{+\infty}\mathrm{e}^{-x\zeta}u_{\mathrm{horz}}(x,y)\,\mathrm{d}x=\mathbf{L}[f](\zeta)-\mathbf{L}[J_{1,0}(y,\cdot)\ast f](\zeta)\\ =\mathbf{L}[f](\zeta)-\mathbf{L}[J_{1,0}(y,\cdot)](\zeta)\mathbf{L}[f](\zeta)=\mathrm{e}^{-y/\zeta}\mathbf{L}[f](\zeta),\qquad\mathrm{Re}\,\zeta>\sigma_{0}.$
@@ 3.1. Estimation of an integral
$\psi(t_{0})=-(1-q)\theta(q\theta/\sigma)^{\frac{q}{1-q}}.$
@@ 4.1. Background on quasianalyticity and analytic quasianalyticity
The concept of analytic quasianalyticity is analogous. If we have a function which is $C^{\infty}$-smooth on the real line, and also possesses a (necessarily unique) holomorphic extension to the upper half-plane, we may still ask when the vanishing of all the Taylor coefficients at a given real point forces the function to vanish everywhere. The natural condition on the (locally) uniform control on the growth of the higher order derivatives remains in force, and the question is asked in terms of the weight sequence controlling the growth. Naturally, the one-sided holomorphic extension property means that less stringent growth control is needed to force the function to vanish everywhere. And when we are in the opposite world of analytic non-quasianalytic classes, that means that that there exist nontrivial functions in the class that are flat at a given real point and holomorphic in the upper half-plane. How should such functions look like? We cannot as in the classical non-quasianalytic case just take such a function and replace it by the 0 function on one side of the given point, as that would ruin the one-sided holomorphic extension property. Typically, if such flat functions exist, there also exist such flat functions with a decay at the given flat point which from the positive imaginary direction is like that of an atomic inner function in the inner-outer factorization theory of the upper half-plane. In [5], Beurling studied weighted spaces on semigroups where analytic non-quasianalyticity is a natural dividing property after application of the Fourier-Laplace transform. Dually, we have instead a space of functions that are defined and holomorphic in the upper half-plane but need not have well-defined boundary values on the real line pointwise. Instead, the functions meet a growth condition in the upper half-plane. Since basically Fourier analysis on the half-line amounts to expansion of the functions in terms of the singular atomic inner functions at the point at infinity, it becomes natural to ask when this carries over to the setting of a finite point in a stable fashion. This stability would require the various atomic inner functions at the given point $x_{0}\in\mathbb{R}$, naturally indexed by a parameter $\alpha\in\mathbb{R}_{\geq 0}$, to generate invariant subspaces $I_{\alpha}(x_{0})$ which get smaller as $\alpha$ increases, and the containments should be strict. This issue appeared later in the work of Korenblum [22] (also in collaboration with Hayman [15]). It was motivated by Korenblum’s earlier work [20], [21] which extended as far as possible the classical Nevanlinna theory as well as the Beurling invariant subspace theorem to the soft topology slow growth spaces $\mathcal{A}^{-\infty}$ consisting of entire functions of finite exponential type in the context of the Poincaré model of the hyperbolic plane. What we will need here is basically the property that the limit of the invariant subspaces $I_{\alpha}(x_{0})$ as $\alpha\to+\infty$ is the trivial subspace $\{0\}$, interpreted in the sense of intersection.
@@ 4.2. A boundary decay uniqueness theorem
can be achieved for a nontrivial bounded analytic function. Note that we can use a conformal transformation (or, more generally, a holomorphic transformation), to achieve a weakening of the assumptions of the theorem. This idea is one of the basic ingredients which helped Korenblum developed his non-Nevanlinna theory in [20], [21]. Here, the more precise setting is in part taken from the paper of Hayman and Korenblum [14], [15].
@@ 5.1. Skewed exponential growth
which defines a holomorphic function in the indicated half-plane for each $y\in\mathbb{R}_{\leq 0}$. In view of (2.6),
@@
${(q\theta_{1})^{\frac{1}{q}}}{(q^{\prime\prime}\theta_{2})^{\frac{1}{q^{\prime\prime}}}}<\cos^{2}\frac{\pi(1-q)}{2q}<1,$
@@
${(q\theta_{1})^{\frac{1}{q}}}{(q^{\prime\prime}\theta_{2})^{\frac{1}{q^{\prime\prime}}}}=\cos^{2}\frac{\pi\alpha}{2},\qquad\frac{1-q}{q}<\alpha<1,$

### 25 Mar 2026 v3

@@ Abstract.
We study the Klein-Gordon equation in one spatial and one temporal dimension. Physically, this equation describes the wave function of a relativistic spinless boson with positive rest mass. Mathematically, this is the most elementary hyperbolic partial differential equation, after the wave equation itself. Relative to the point at the origin, the spacetime splits according to the light cones, and we find four quarter-planes, two of which are timelike while the remaining two are spacelike. Not unexpectedly, the solutions behave quite differently in the two types of quarter-planes. It turns out that the spacelike quarter-planes exhibit a Liouville phenomenon, where insufficient growth forces the solution to display a certain kind of symmetry, where the values along the two linear edges are in a one-to-one relation. This phenomenon shares features with the classical Liouville theorem in the theory of entire functions as well as the Phragmén-Lindelöf principle for harmonic functions.
@@ 1.1. Classical backgrund
The Klein-Gordon equation describes the wavefunction of a relativistic spinless boson with positive rest mass. In one spacial and one temporal dimension, it is advantageous to swith to using the characteristic directions to define alternative coordinates, which after rescaling gives us the equation
@@
since $f(0)=g(0)$, $J_{0,0}(x,0)=1$, and $J_{0,1}(x-s,0)=0$. By interchanging the roles of $x$ and $y$ in the calculation, we find likewise that
@@
To offer a cohesive reader-friendly presentation, we supply the proof of this theorem in the appendix (Section 7). Riemann’s contribution to the Darboux-Goursat problem with data along characteristics can be found in his collected works [24] (the paper Ueber die Fortplantzung ebener Luftwellen von endlicher Schwingungsweite is the most relevant), and for this reason, the method is called the Riemann method.
@@
If we apply (1.6) to $a^{\prime}=x,b^{\prime}=y,a=b=0$, we may view the equation as a way to represent $u(x,y)$ in terms of the values $u(x,0)$, $u(0,y)$, $u(0,0)$, together with a Volterra-type integral:
@@
Based on this representation, the Picard method applies to give uniqueness.
@@ 1.5. Splitting Riemann’s solution into unilateral wave solutions
This assertion is immediate from Theorems 1.2 and 1.3 by slightly rearranging the terms in Riemann’s solution. Alternatively, we can observe that the expressed solution $u$ solves the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ in the sense of distribution theory since each of the terms $u_{0},u_{\mathrm{horz}},u_{\mathrm{vert}}$ does, and that the values along the the $x$-axis and the $y$-axis coincide with $f$ and $g$, respectively. Theorem 1.3 then shows that $u=u_{0}+u_{\mathrm{horz}}+u_{\mathrm{vert}}$ must coincide with $\mathbf{R}[f,g]$. Proposition 1.8 motivates the following terminology.
@@
It is the aim of this paper to study the growth properties of the unilateral wave solutions $u_{\mathrm{horz}}$ and $u_{\mathrm{vert}}$. By Lorentz group invariance, it is equivalent to focus on just one of them, and we will consider the horizontal unilateral wave solution $u_{\mathrm{horz}}$. As it turns out, there is a Liouville phenomenon for such solutions. Here, we recall the classical theorem of Liouville, which asserts that any entire function which is also bounded has to be constant. There are many extensions, for instance, a harmonic function in the entire plane which is bounded must be constant as well. A version which is particularly attractive is known as the Phragmén-Lindelöf principle which asserts that if a harmonic function vanishes on a line and has insufficient growth on one side of the line, then it must vanish everywhere.
@@ 1.6. Motivation: the Phragmén-Lindelöf principle for harmonic functions
The connection with Liouville’s theorem is actually rather direct: By Schwarz reflection, which applies along the linear boundary since the values of the harmonic function vanish there, the harmonic function $u$ extends to a harmonic function in the whole complex plane with $u(\bar{z})=-u(z)$, where we identify $z=x+\mathrm{i}y\sim(x,y)$. The growth assumptions on this entire harmonic function are such that we have very slow growth. In other words, we are close in spirit to the classical Liouville theorem.
@@
There is a rather sizable literature on Liouville type phenomena for elliptic equations, linear and nonlinear, also involving fractional Laplacians, etc. We could mention a particularly interesting effect in the discrete context, where harmonicity combined with boundedness on a big enough portion of the lattice $\mathbb{Z}^{2}$ is enough to force the function to be constant [7]. This is in contrast with the continuous setting where it is enough to have a narrow path tending to infinity to permit the existence of a nontrivial harmonic function which grows wildly along the path while it is bounded elsewhere.
@@ 1.7. Liouville type phenomenon for spacelike quarter-planes
Apparently, in the context of the Phragmén-Lindelöf principle, there is some degree of minimal growth which must be surpassed to get a nontrivial solution. What about the unilateral solutions to the Klein-Gordon equation? By the characterization of Proposition 1.10, they vanish either on the horizontal line $\mathbb{R}\times\{0\}$ or the vertical line $\{0\}\times\mathbb{R}$. This may be able to force them to grow somewhere, and, in fact, as it turns out, the required growth is rather strong. A first result of this kind which we may present is the following. Note that with respect to the origin $(0,0)$, the equation $xy=0$ represents the light cone emanating from $(0,0)$, while $xy>0$ represents the timelike vectors, and $xy<0$ represents the spacelike vectors. The spacelike vectors can all be thought of as representing events contemporary with that of the origin, and hence without causal relationship to the origin.
@@
The proof of Theorem 1.14 is supplied in Subsection 2.3. A reasonable interpretation of the theorem is that for solutions relevant to Physics, the values along the negative $y$-axis determine the values along the positive $x$-axis, and vice versa.
@@
(c) In the instance that $\mathcal{Y}=\mathbb{R}_{\leq 0}$, it appears that Theorem 1.14 was first obtained by Andrew Bakan (unpublished). After all, he mentioned a version of it in a seminar presentation in Stockholm during his visit in the fall semester of 2016. We have since learned that Bakan’s interest in the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ has its roots in the 1990s. However, the paper by Peter Ullrich [25] also gives unique extension of solutions to the Klein-Gordon equation from a characteristic hypersurface in a higher-dimensional context, where the growth control is not pushed to the limit (basically polynomial growth control, as this corresponds to the growth of tempered distributions). We can also refer to Heinzl and Werner for some related physical considerations [20].
The condition that $u$ vanishes along the semi-axis $\{0\}\times\mathbb{R}_{\leq 0}$ entails that $u$ may be extended to a solution to the Klein-Gordon equation in the whole plane $\mathbb{R}^{2}$ which vanishes on the line $\{0\}\times\mathbb{R}$, that is, the extension is a horizontal unilateral solution.
@@
Suppose $\theta_{1},\theta_{2}\in\mathbb{R}_{>0}$ with $\theta_{1}\theta_{2}\geq 1$. There exists a nontrivial continuous horizontal unilateral solution $u=u_{\mathrm{horz}}$ to the Klein-Gordon equation $u^{\prime\prime}_{xy}+u=0$ such that the restriction of $u$ to the quarter-plane $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$ does not vanish identically, while it meets the uniform growth condition
@@
$\theta_{1}^{1/q}\theta_{2}^{1/q^{\prime\prime}}<\bigg(\frac{1}{q}\bigg)^{1/q}\bigg(\frac{1}{q^{\prime\prime}}\bigg)^{1/q^{\prime\prime}}\sin^{2}\frac{\pi}{2q},$
@@
$\theta_{1}^{1/q}\theta_{2}^{1/q^{\prime\prime}}>\bigg(\frac{1}{q}\bigg)^{1/q}\bigg(\frac{1}{q^{\prime\prime}}\bigg)^{1/q^{\prime\prime}}\sin^{2}\frac{\pi}{2q},$
@@ 2.1. An example of growth in a spacelike quadrant
holds uniformly. Moreover, since $u_{1}(x,0)=f_{1}(x)$ holds for $x\in\mathbb{R}$, it follows that $u_{1}$ is a nontrivial solution to the Klein-Gordon equation which in particular does not vanish identically on $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\leq 0}$ while it enjoys the claimed growth estimate. ∎
@@ 2.2. The evolution equation for horizontal unilateral wave solutions
(2.6) $\mathbf{L}[u_{\mathrm{horz}}(\cdot,y)](\zeta)=\int_{0}^{+\infty}\mathrm{e}^{-x\zeta}u_{\mathrm{horz}}(x,y)\,\mathrm{d}x=\mathbf{L}[f](\zeta)-\mathbf{L}[J_{0,1}(\cdot,y)\ast f](\zeta)\\ =\mathbf{L}[f](\zeta)-\mathbf{L}[J_{0,1}(\cdot,y)](\zeta)\mathbf{L}[f](\zeta)=\mathrm{e}^{-y/\zeta}\mathbf{L}[f](\zeta),\qquad\mathrm{Re}\,\zeta>\sigma_{0}.$
@@ 3.1. Estimation of an integral
$\psi(t_{0})=-\theta(1-q)(q\theta/\sigma)^{\frac{q}{1-q}}.$
@@ 4.1. Background on quasianalyticity and analytic quasianalyticity
The concept of analytic quasianalyticity is analogous. If we have a function which is $C^{\infty}$-smooth on the real line, and also possesses a (necessarily unique) holomorphic extension to the upper half-plane, we may still ask when the vanishing of all the Taylor coefficients at a given real point forces the function to vanish everywhere. The natural condition on the (locally) uniform control on the growth of the higher order derivatives remains in force, and the question is asked in terms of the weight sequence controlling the growth. It is to be expected that the one-sided holomorphic extension property means that less stringent growth control is needed to force the function to vanish everywhere. And when we are in the opposite world of analytic non-quasianalytic classes, that means that that there exist nontrivial functions in the class that are flat at a given real point and holomorphic in the upper half-plane. What should such functions look like? We cannot as in the classical non-quasianalytic case just take such a function and replace it by the 0 function on one side of the given point, as that would ruin the one-sided holomorphic extension property. Typically, if such flat functions exist, there also exist such flat functions with a decay at the given flat point which from the positive imaginary direction is like that of an atomic inner function in the inner-outer factorization theory of the upper half-plane. In [4], Beurling studied weighted spaces on semigroups where analytic non-quasianalyticity is a natural dividing property after application of the Fourier-Laplace transform. Dually, we have instead a space of functions that are defined and holomorphic in the upper half-plane but need not have well-defined boundary values on the real line pointwise. Instead, the functions meet a growth condition in the upper half-plane. Since basically Fourier analysis on the half-line amounts to expansion of the functions in terms of the singular atomic inner functions at the point at infinity, it becomes natural to ask when this carries over to the setting of a finite point in a stable fashion. This stability would require the various atomic inner functions at the given point $x_{0}\in\mathbb{R}$, naturally indexed by a parameter $\alpha\in\mathbb{R}_{\geq 0}$, to generate invariant subspaces $I_{\alpha}(x_{0})$ which get smaller as $\alpha$ increases, and the containments should be strict. This issue appeared later in the work of Korenblum [23] (also in collaboration with Hayman [14]). It was motivated by Korenblum’s earlier work [21], [22] which extended as far as possible the classical Nevanlinna theory as well as the Beurling invariant subspace theorem to the soft topology slow growth spaces $\mathcal{A}^{-\infty}$ consisting of entire functions of finite exponential type in the context of the Poincaré model of the hyperbolic plane. What we will need here is basically the property that the limit of the invariant subspaces $I_{\alpha}(x_{0})$ as $\alpha\to+\infty$ is the trivial subspace $\{0\}$, interpreted in the sense of intersection.
@@ 4.2. A boundary decay uniqueness theorem
can be achieved for a nontrivial bounded analytic function. Note that we can use a conformal transformation (or, more generally, a holomorphic transformation), to achieve a weakening of the assumptions of the theorem. This idea is one of the basic ingredients which helped Korenblum to develop his non-Nevanlinna theory in [21], [22]. Here, the more precise setting is in part taken from the paper of Hayman and Korenblum [13], [14].
@@ 5.1. Skewed exponential growth
which defines a holomorphic function in the indicated half-plane for each $y\in\mathcal{Y}$. In view of (2.6),
@@
${(q\theta_{1})^{\frac{1}{q}}}{(q^{\prime\prime}\theta_{2})^{\frac{1}{q^{\prime\prime}}}}<\sin^{2}\frac{\pi}{2q}=\cos^{2}\frac{\pi(1-q)}{2q}<1,$
@@
${(q\theta_{1})^{\frac{1}{q}}}{(q^{\prime\prime}\theta_{2})^{\frac{1}{q^{\prime\prime}}}}=\cos^{2}\frac{\pi\alpha}{2},\quad\text{where}\quad\frac{1-q}{q}<\alpha<1,$